OpenStudy (anonymous):

Pedro has created the function f(x) = the quantity of 4x minus 3, divided by 2 to represent the number of assignments he has completed, where x represents the number of weeks in the course. Pedro discovers that using the inverse function to solve for x = 30, he can predict when he will have 30 assignments completed. Explain to Pedro how to accomplish this, using complete sentences. @helpme1.2

OpenStudy (anonymous):

f(x)= (4x-3)/2 30 = (4x-3)/2 solve for x

OpenStudy (anonymous):

how :(

OpenStudy (anonymous):

30 = (4x-3)/2 -- cross multiply 60 = 4x-3 --- add 3 to both sides 63 = 4x -- divide both by 4 15.75=x

OpenStudy (anonymous):

you can explain that in a complete sentence now